OpenStudy (anonymous):

^3√(3x-1)=2 show all work please

OpenStudy (anonymous):

everything in ( ) is under the square root

geerky42 (geerky42):

Cube both sides.

OpenStudy (johnweldon1993):

Ahh okay that clears the last question up \[\sqrt[3]{3x-1} = 2\] you would cube both sides to cancel out the cubed root and you're left with 3x - 1 = 2^3 I believe you should be able to do this now?

OpenStudy (anonymous):

Would x=8 ?]

geerky42 (geerky42):

None. You are not done. After cube the equation, you have 3x - 1 = 2³ You have to solve for x now. 3x - 1 = 2³ 3x - 1 = 8 3x = 9 \(\vdots\) is this clear?

OpenStudy (anonymous):

Yes very Clear ! Thank you so much.